OpenStudy (anonymous):

If y varies directly with x and y = -7 when x = -1, find x when y = -84. Answer 84 7 -12

OpenStudy (phi):

can you find k ? in y = k x

OpenStudy (anonymous):

how do you find k

OpenStudy (phi):

use y = -7 when x = -1, in y = k x that means replace y with -7 and x with -1 what do you get ?

OpenStudy (jdoe0001):

same as before http://openstudy.com/study#/updates/528fe8a1e4b0424725722e9d

OpenStudy (anonymous):

-7=k*-1

OpenStudy (phi):

now divide both sides by -1. on the right side -1 divided by -1 is 1

OpenStudy (phi):

\[ \frac{-7}{-1}= \frac{-1}{-1} k \]

OpenStudy (anonymous):

ok so -7=k*1

OpenStudy (phi):

almost. -7 divided by -1 is +7 (two minus signs make a positive)

OpenStudy (anonymous):

oh my calculator didnt take away the negetive ok i get it now

OpenStudy (phi):

you now know k= 7 so your equation y= k x is now y = 7 x now if you want y for some x, replace x with the number, and multiply it by 7

OpenStudy (phi):

or in this case, replace y with -84

OpenStudy (anonymous):

i am confused

OpenStudy (phi):

do you understand how we got y = 7 x ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the answer is 7!

OpenStudy (phi):

they say y = -84. in y = 7 x replace y with -84 -84 = 7 x to find x, divide both sides by 7 -84/7 = x I don't think that is 7